The length of a rectangle is twice its width and its perimeter is 120 cm. Find the dimensions of the rectangle. I know the answer is 40cm x20cm but i need the formula
Answer by Earlsdon(6103)

L = 2W and...
P = 120, so...
2(L+W) = 120 Substitute L = 2W.
2((2W)+W) = 120 Divide both sides by 2 and simplify the left side.
3W = 60 Divide both sides by 3.
W = 20 and L = 2(20) = 40